### v4

# LLM-Based Emulation of the Radio Resource Control Layer: Towards AI-Native RAN Protocols

Abstract
Integrating Large AI Models (LAMs) into 6G mobile networks is a key enabler of the AI‑Native Air Interface (AI-AI), where protocol intelligence must scale beyond handcrafted logic. This paper presents, to our knowledge, the first standards-compliant emulation of the Radio Resource Control (RRC) layer using a decoder-only LAM (LLaMA-class) fine-tuned with Low‑Rank Adaptation (LoRA) on a multi-vendor corpus of real-world traces spanning both 5G and 4G systems. We treat RRC as a domain-specific language and construct a segmentation-safe, question–answer (Question-and-Answer (QA)) dataset that preserves Abstract Syntax Notation (ASN.1) structure through linearization prior to Byte Pair Encoding (BPE) tokenization. The proposed approach combines parameter-efficient adaptation with schema-bounded prompting to ensure syntactic and procedural fidelity. Evaluation introduces a standards-aware triad—ASN.1 conformance, field-level coverage analysis, and uplink-to-downlink state-machine checks—alongside semantic similarity and latency profiling across 120 configurations. On 30k 5G request–response pairs plus an additional 4.8k QA turns from 4G sessions, our 8B model achieves a median cosine similarity of 0.97, a 61% relative gain over a zero-shot baseline, while sustaining high conformance rates. These results demonstrate that LAMs, when augmented with protocol-aware reasoning, can directly orchestrate control-plane procedures, laying the foundation for the future Artificial Intelligence (AI)-native Radio Access Network (RAN).
I Introduction
Large AI Models (LAMs) are poised to transform wireless networks by enabling autonomous, cognitive capabilities that go beyond the scope of traditional optimization methods[3]. The forthcoming sixth generation (6G) of mobile networks is expected to incorporate Artificial Intelligence (AI) methods not merely as auxiliary optimisation tools but also as integral design primitives of the AI‑Native Air Interface (AI-AI). Recent proposals for an AI-AI posit that Machine Learning (ML) models could co‑design Physical Layer (PHY) and Medium Access Control (MAC) procedures [15], thereby enabling radio stacks that adapt automatically to dynamic service requirements, resource constraints, and propagation conditions [4]. The rapid evolution of LAMs [10, 13] strengthens this vision—when endowed with multi-modal perception and reasoning, such models become plausible agents for protocol‑centric tasks that demand both linguistic competence and structured decision making. This raises a timely research challenge: determining the extent to which LAMs can directly engage in the generation and interpretation of control signaling in the Radio Access Network (RAN). Can LAMs achieve sufficient protocol literacy for RRC [1] emulation? What fine-tuning strategies are effective? What is the performance fidelity of such a system? These are the research questions this paper intends to answer. We hypothesize that LAM‑scale pretraining provides the compositional priors needed to generalize across 3GPP releases and vendor variants, and to perform multi‑procedure, stateful reasoning under incomplete context—capabilities that compact task‑specific transformers or grammar‑only decoders lack without extensive bespoke engineering.
One particularly compelling target for AI intervention is the Radio Resource Control (RRC) layer of 5G New Radio (NR). RRC is the Layer 3 (L3) of the cellular network. It orchestrates connection establishment, mobility and configuration between User Equipment (UE) and the Next‑Generation Node B (gNB). Conceptually, RRC messages form a specialized, domain‑specific language; learning this language is analogous to learning syntax and semantics in natural‑language processing. Treating RRC messages as such positions LAM ‑based sequence models as natural emulators of the RRC layer. Demonstrating accurate RRC emulation would constitute an early, concrete instantiation of AI-AI principles, with direct implications for the design of the 6G control plane.
Can the protocol literacy of a LAM —specifically, its ability to parse and generate 3GPP-compliant RRC messages—be systematically improved, and through which fine‑tuning strategies? This paper answers this question through a novel two‑phase methodology that (i) imparts RRC knowledge to a decoder‑only LAM via supervised fine‑tuning enhanced with Low‑Rank Adaptation (LoRA), and (ii) evaluates the resulting model’s fidelity in emulating RRC procedures under realistic network scenarios. The findings offer evidence about the feasibility of LAM ‑based standards-compliant RRC emulation.
Without loss of generality, this work focuses on decoder‑only transformer architectures (e.g., Llama‑class models [14]). Autoregressive decoding matches the turn‑taking structure of RRC exchanges and permits key–value caching, which curtails inference latency and memory footprint, both critical for deployment at the gNB. Decoder‑only models thereby strike a favourable balance between capacity and computational efficiency, especially when combined with parameter‑efficient fine‑tuning techniques such as LoRA [5]. LoRA freezes the original model weights and introduces low‑rank update matrices, reducing trainable parameters by orders of magnitude while maintaining expressiveness; the approach mitigates both the cost of domain adaptation and the runtime overhead of inference, two primary barriers to operational use.
Training LAMs for specialist telecom domains demands curated datasets that capture the combinatorial breadth of protocol states, while deployment must meet stringent latency budgets to avoid impairing control‑plane responsiveness. To address these issues, we rebuilt two data corpuses of RRC traces (4G and 5G) with multi-operator coverage. Then, we adopt a standards-aware evaluation suite that combines ASN.1 [6] syntax validation, schema-level field coverage, and UL→DL state-machine conformance, reporting these jointly with semantic similarity and latency across 120 configurations. Ultimately, we also outline a forward-looking engineering plan toward sub-100 ms inference latency via quantization [2], constrained decoding, KV caching and multi‑token draft‑verify decoding.
II System Model
This section delineates the architectural and operational principles underlying the proposed LAM-integrated gNB. We first describe the system architecture, detailing the functional decomposition of the disaggregated gNB and the embedding of a LAM within the RRC layer, while ensuring adherence to standardized interface definitions. Then, the model-driven control-plane workflow is examined, encompassing the reception and contextual interpretation of uplink RRC messages and the autoregressive synthesis of standards-compliant downlink responses.
Figure 1: High-level concept of an AI-native gNB-side RRC layer powered by a LLM, illustrating its core inputs, outputs, learning source, and interaction with the protocol stack.
II-A Architecture and Functional Integration
The proposed Large RRC Model integrates a decoder-only transformer-based LAM into the RRC layer of a disaggregated gNB, as depicted in Fig. 1. Following the 3rd Generation Partnership Project (3GPP) NG-RAN standard, the gNB is partitioned into a Central Unit (CU) and a Distributed Unit (DU), with the CU further subdivided into separate Control Plane of Central Unit (CU-CP) and User Plane of Central Unit (CU-UP) entities. The CU-CP hosts the RRC layer, responsible for essential control-plane tasks such as connection establishment, radio bearer management, and mobility decisions. Replacing the traditional rule-based logic, the embedded RRC LLM serves as an intelligent decision engine within the RRC, producing control messages and associated protocol decisions via autoregressive inference. Conversely, the CU-UP encompasses user-plane layers, primarily Service Data Adaptation Protocol (SDAP) and user-plane Packet Data Convergence Protocol (PDCP), that manage the transport of user data. The DU, hosting lower-layer (e.g., Radio Link Control (RLC), MAC, and PHY layers) protocols, interfaces directly with the Remote Radio Heads (RRHs) performing Radio Frequency (RF) and low-level physical-layer tasks. Fig. 2 illustrates all these interfaces and functions. Integrating an LLM at the RRC level is intended to preserve the external protocol behavior, leaving northbound and lateral interfaces to peer nodes unaltered, thus ensuring interoperability with existing network deployments.
Figure 2: Low‑level architecture of an NR gNB disaggregating an LLM‑based RRC layer.
As in standard 5G New Radio (NR), uplink RRC messages from UEs, such as Connection Requests or Measurement Reports, arrive at the CU-CP via Control‑Plane Interface between CU‑CP and DU (F1-C), where they are processed by the embedded RRC LLM, replacing the conventional RRC layer implementation. The model interprets these inputs contextually, considering ongoing connection states, historical interactions, and dynamic network conditions to generate suitable downlink RRC responses (e.g., Connection Setup or Mobility Reconfiguration). The generated messages maintain compliance with standardized 3GPP encoding, thus transparently replacing the message generation logic. Downlink responses are then transmitted back to the UE through the DU via the same interface.
II-B LLM-Driven Signaling Workflow and Model Operation
Interactions between the RRC layer and the user-plane modules are similarly orchestrated. When the Large RRC Model decides to initiate or modify data radio bearers—potentially based on inferred context or changing service demands—it signals these decisions to the CU-UP over the Interface between CU-CP and CU-UP (E1). The CU-UP subsequently configures the corresponding SDAP and PDCP entities, enabling adaptive management of the user-plane traffic. Meanwhile, data-plane traffic flows remain unmodified and proceed independently over the established user-plane interfaces (User‑Plane Interface between CU‑UP and DU (F1-U), User‑Plane Interface between RRH and DU (F2-U)).
Figure 3: Data flow for fine-tuning and inference in the LLM-based RRC system.
Fig. 3 illustrates the fine-tuning and inference workflow specific to the proposed RRC LLM. The decoder-only transformer is taught RRC domain knowledge using supervised fine-tuning on a curated dataset of historical RRC message traces collected from a real-world 5G network. This specialized fine-tuning imparts explicit protocol knowledge to the model, enabling it to emulate standard RRC behaviors accurately during live inference. Operationally, the trained model functions continuously within the CU-CP, dynamically generating control-plane messages as new UE requests are received.
III Fine-tuning of LLM for RRC Message Processing
This section describes the pipeline for adapting a decoder-only LLM to emulate RRC procedures. Our approach consists of three components: (i) dataset construction, where raw traces are reorganized into segmentation-safe uplink–downlink Question-and-Answer (QA) pairs; (ii) parameter-efficient supervised fine-tuning using LoRA; and (iii) latency-aware evaluation to assess feasibility under control-plane timing constraints. Earlier traces often exceeded context limits, requiring ad-hoc segmentation that broke procedural continuity. Our design eliminates this by enforcing per-procedure boundaries.
III-A Datasets: Construction, Scope, and Coverage
We use two complementary corpora with a common, segmentation-safe construction pipeline:
[…]
  • LTE dataset (proprietary road tests across two networks): $\sim$4,808 QA turns across 1,601 sessions collected via real-world drive tests.
[…]
Raw RRC traces are sessionized and reorganized into UL–DL QA pairs preserving procedural causality and ASN.1 structure (cf. Figure 4). Each session is capped in turns to respect context limits, with adjacent UL messages concatenated when appropriate (e.g., rrcReconfigurationComplete ++ measurementReport). The resulting per-turn inputs ($X$) and targets ($Y$) follow the mapping in section III-B and Eq. 14.
Schema-bounded prompting and radio constraints
To enforce structural validity, we inject a session-specific, size-bounded ASN.1 micro-schema (see Fig. A.1 in the appendix) into the system prompt together with radio constraints (e.g., allowed EARFCNs). This preserves field names/order and prevents extraneous elements without token-level grammar overhead. This micro-schema mechanism was applied only to the LTE corpora.
Attribution and availability
Both corpora are private and proprietary. Public LTE datasets/simulators exist (e.g., [9, 8]), but are synthetic and were not used in this work.
Coverage
Table I summarizes the observed message-type counts across the two corpora. Major RRC procedures (connection establishment, initial security activation, reconfiguration, release, capability transfer) are well represented in both RATs.
TABLE I: Breakdown of LTE and NR RRC message types observed in the proprietary field logs. A — denotes not observed for that RAT.
Message LTE Count NR Count
dlInformationTransfer 149 1 743
measurementReport 229 43 410
mobilityFromNRCommand 92
paging 1 016
rrcConnectionReconfiguration 701
rrcConnectionReconfigurationComplete 891
rrcConnectionRelease 110
rrcConnectionRequest 662
rrcConnectionReestablishment 3
rrcConnectionReestablishmentRequest 5
rrcConnectionSetup 775
rrcConnectionSetupComplete 879
rrcReconfiguration 16 034
rrcReconfigurationComplete 16 099
rrcRelease 4 267
rrcReestablishment 28
rrcReestablishmentComplete 28
rrcReestablishmentRequest 119
rrcSetup 4 072
rrcSetupComplete 4 068
rrcSetupRequest 4 009
securityModeCommand 653 3 827
securityModeComplete 693 3 824
systemInformation 95
systemInformationBlockType1 265
ueCapabilityEnquiry 633 876
ueCapabilityInformation 6 873
ueInformationRequest-r9 3
ueInformationResponse-r9 134
ulInformationTransfer 55 1 924
Total 7 957 105 293
[…]
[width=]figures_latex/rrc_trace
Figure 4: Example of a pre-processed RRC trace segment. Messages are structured into Question (Q) / Answer (A) pairs for LLM training, where ’Q’ denotes the input RRC message(s) (X) and ’A’ the target response (Y). Note the consolidation of multiple physical messages (e.g., rrcReconfigurationComplete and measurementReport) into a single logical request, labeled here as Q4(concat) and Q6(concat).
III-B Supervised Fine-Tuning Framework
We model RRC request–response exchanges as a conditional generation task over uplink–downlink pairs. Let $X$ denote the uplink message sequence and $Y$ the corresponding downlink response. The objective is to estimate the conditional probability:
$\pi_{\theta}(Y\mid X)=\prod_{t=1}^{T}P(y_{t}\mid y_{<t},X;\theta),$ (1)
where $\theta$ are trainable parameters, $y_{t}$ is the token generated at step $t$, and $T$ is the sequence length. This formulation aligns with the QA-style mapping introduced in Sec. III-A.
To adapt a decoder-only LLM efficiently, we apply LoRA, which introduces low-rank updates to frozen backbone weights. For a weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the update is:
$W_{\theta}=W_{0}+AB,$ (2)
where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$ are trainable factors with rank $r\ll\min\{d,k\}$. This reduces trainable parameters by orders of magnitude while preserving expressiveness. The fine-tuning objective minimizes the negative log-likelihood:
$\mathcal{L}_{\text{SFT}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log\pi_{\theta}(Y^{(i)}\mid X^{(i)}).$ (3)
Here, $N$ denotes the total number of aligned uplink–downlink pairs in the supervised training set. Inference uses greedy decoding under the fine-tuned model:
$\hat{Y}=\arg\max_{Y}\pi_{\theta}(Y\mid X).$ (4)
Algorithm 1 RRC-LLM Training and Inference Pipeline
1:Raw RRC traces $\mathcal{T}_{\text{raw}}$, pre-trained model $\pi_{\theta_{\text{init}}}$
2:Fine-tuned model $\pi_{\theta^{*}}$
3:Phase 1: Dataset Construction
[…]
5:for all $(X^{(i)},Y^{(i)})\in\mathcal{D}$ do
6:  $X^{(i)}\leftarrow\texttt{BPE\_Tokenize}(X^{(i)})$
7:  $Y^{(i)}\leftarrow\texttt{BPE\_Tokenize}(Y^{(i)})$
8:end for
9:Phase 2: Supervised Fine-Tuning (SFT)
[…]
11:for $e=1$ to $E_{\text{SFT}}$ do
12:  for all $(X^{(i)},Y^{(i)})\in\mathcal{D}$ do
13:   $\mathcal{L}_{\text{SFT}}\leftarrow-\log\pi_{\theta_{\text{ref}}}(Y^{(i)}\mid X^{(i)})$
14:   Update $\theta_{\text{ref}}$ using Adam on $\mathcal{L}_{\text{SFT}}$
15:  end for
16:end for
17:Phase 3: Inference
18:Given new uplink message $X$, compute
19:   $\hat{Y}\leftarrow\arg\max_{Y}\pi_{\theta}(Y\mid X)$ return $\hat{Y}$
III-C Training Configuration and Convergence Behavior
Supervised fine-tuning of the Llama-3 8B backbone was performed under the configuration in Table II. The loss decreased by two orders of magnitude within the first $10^{3}$ steps and stabilized after $\sim$8.4k steps, as shown in Fig. 5. The EMA-smoothed trace flattens near $10^{-2}$, indicating convergence. We select the checkpoint at this point to capture $95\%$ of attainable loss reduction while avoiding overfitting and unnecessary GPU hours. Fig. 6 further compares training and validation loss across backbones and LoRA ranks, highlighting stability trends and the effect of parameter‑efficient tuning.
[…]
Figure 5: Training loss trajectory on the NR dataset (proprietary corpus). The pale blue trace is the raw loss, and the dark blue curve is its exponential moving average (EMA). The vertical red dashed line marks the epoch selected for downstream evaluation.
TABLE II: Supervised fine-tuning hyperparameters
Parameter NR Dataset LTE Dataset
Model Llama-3 8B Llama-3.x (1B–8B)
Max Sequence Length 4096 4096
Precision bf16 bf16
Optimizer AdamW AdamW
Learning Rate $2\times 10^{-5}$ $2\times 10^{-5}$ (full FT) $3\times 10^{-4}$ (LoRA)
Loss Function CrossEntropy CrossEntropy
LR schedule Cosine decay + 100-step warmup
LoRA Config $\alpha=2r$, dropout=0
III-D Decoding regimes
We evaluate three decoding strategies that govern prompt structure and output constraints:
  • NoSys: A generic system prompt with minimal guidance, allowing unconstrained generation. This regime prioritizes raw language-model priors and serves as a lower-bound baseline.
  • RRC: A domain-oriented prompt that specifies the RRC context and message type but does not enforce strict schema constraints. This improves semantic alignment but leaves structural validity largely unchecked.
  • RRC_constrain: A schema-bounded prompt that injects a size-limited Abstract Syntax Notation (ASN.1) micro-schema and radio constraints (e.g., allowed E-UTRA Absolute Radio Frequency Channel Number (EARFCN)) into the system message. Generation is restricted to fields declared in this schema, preserving field order and prohibiting extraneous elements. This regime is designed to maximize standards compliance without incurring token-level grammar overhead.
Unless otherwise stated, all reported results in Sec. IV use RRC_constrain for tuned configurations, as it delivers near-ceiling ASN.1/State-Machine Conformance (SMC) pass rates while maintaining practical latency.
III-E LTE Corpus: Multi-Backbone and LoRA Rank Ablations
To evaluate cross-RAT robustness, we repeated the fine-tuning pipeline on an LTE-specific corpus (1,601 sessions, 4,808 QA turns) using the same segmentation-safe design as in Sec. III-A. This LTE dataset captures connection setup, security activation, and reconfiguration procedures under diverse mobility scenarios.
We benchmarked four Llama-3 backbones (8B, 8B-3.1, 3B, 1B) under full fine-tuning and LoRA with ranks $r\in\{4,8,16\}$. All runs used bf16 precision, AdamW optimizer, and token-level cross-entropy loss. LoRA adapters were applied to attention projections and MLP layers with scaling $\alpha=2r$ and dropout =0=0. Learning rates followed standard practice: $2\times 10^{-5}$ for full FT and $3\times 10^{-4}$ (cosine decay, 100-step warmup) for LoRA. 5
Table III summarizes the best validation loss per configuration. On 8B backbones, higher-rank LoRA ($r=16$) matches or slightly surpasses full fine-tuning, while on 3B/1B models, updating all weights remains advantageous. Across all settings, the gap between the best two methods is small ($\Delta\approx 6\times 10^{-4}$), confirming that parameter-efficient adaptation can approach full FT performance on large models.
TABLE III: LTE Corpus: Validation Loss at Best Checkpoint (Lower is Better)
Backbone Full FT LoRA r4 LoRA r8 LoRA r16
Llama-3 8B 0.02017 0.02038 0.02062 0.01984
Llama-3.1 8B 0.02836 0.02131 0.02152 0.02096
Llama-3.2 3B 0.01915 0.02125 0.02232 0.02207
Llama-3.2 1B 0.02017 0.02130 0.02106 0.02105
These results highlight two trends: (i) larger backbones benefit from higher LoRA ranks, reducing the need for full FT; (ii) smaller models favor full updates for best accuracy. This informs the latency–fidelity trade-offs explored in Sec. IV.
TABLE IV: Curated results on the LTE dataset.
Backbone Quant Strategy Tuning Latency (ms) Schema Similarity Pass Rate
Med. Avg. Avg Med Avg Med ASN SMC
L-3 8B FP16 RRC_constrain lora-r16 3544 5258.0 0.976 1.000 0.992 1.000 0.995 0.993
L-3 8B Q4_K_M RRC_constrain lora-r16 2556 3854.1 0.969 1.000 0.990 1.000 0.989 0.988
L-3 8B FP16 RRC lora-r16 3418 4280.4 0.467 0.333 0.723 0.669 0.023 0.000
L-3 8B Q4_K_M RRC lora-r16 2519 2529.7 0.439 0.333 0.726 0.666 0.003 0.002
L-3 8B FP16 NoSys lora-r16 1987 2011.7 0.231 0.241 0.480 0.478 0.003 0.002
L-3 8B Q4_K_M NoSys lora-r16 1431 1442.1 0.235 0.239 0.491 0.490 0.012 0.004
L-3.1 8B FP16 RRC_constrain lora-r16 3609 5367.3 0.969 1.000 0.991 1.000 0.995 0.994
L-3.1 8B Q4_K_M RRC_constrain lora-r16 2581 3939.0 0.964 0.987 0.989 1.000 0.904 0.902
L-3.1 8B FP16 RRC lora-r16 3364 2716.1 0.523 0.675 0.643 0.650 0.001 0.000
L-3.1 8B Q4_K_M RRC lora-r16 1379 1653.6 0.371 0.262 0.579 0.508 0.001 0.000
L-3.1 8B FP16 NoSys lora-r16 2193 2243.0 0.211 0.224 0.477 0.478 0.003 0.002
L-3.1 8B Q4_K_M NoSys lora-r16 1501 1537.8 0.213 0.220 0.485 0.482 0.001 0.000
L-3.2 3B FP16 RRC_constrain full 2341 3501.9 0.969 1.000 0.991 1.000 0.993 0.991
L-3.2 3B Q4_K_M RRC_constrain full 1733 2658.1 0.954 1.000 0.987 1.000 0.973 0.968
L-3.2 3B FP16 RRC full 2157 2939.1 0.408 0.353 0.688 0.700 0.000 0.000
L-3.2 3B Q4_K_M RRC full 2170 2920.6 0.440 0.400 0.689 0.698 0.001 0.000
L-3.2 3B FP16 NoSys full 1755 1848.0 0.214 0.223 0.466 0.467 0.011 0.004
L-3.2 3B Q4_K_M NoSys full 1367 1452.8 0.201 0.207 0.497 0.493 0.040 0.015
L-3.2 1B FP16 RRC_constrain full 2093 3867.2 0.964 1.000 0.988 1.000 0.993 0.989
L-3.2 1B Q4_K_M RRC_constrain full 2025 3585.8 0.934 0.987 0.981 1.000 0.888 0.883
L-3.2 1B FP16 RRC full 589 1251.0 0.336 0.333 0.652 0.649 0.000 0.000
L-3.2 1B Q4_K_M RRC full 428 1368.8 0.293 0.286 0.662 0.665 0.000 0.000
L-3.2 1B FP16 NoSys full 1142 1390.6 0.202 0.200 0.521 0.514 0.024 0.009
L-3.2 1B Q4_K_M NoSys full 1209 1335.5 0.189 0.182 0.501 0.502 0.024 0.016
Figure 6: Training vs. validation loss across models and methods on the LTE dataset (public corpus). Columns: full fine‑tuning and LoRA with $r\!\in\!\{4,8,16\}$. Rows: Llama‑3 8B, Llama‑3.1 8B, Llama‑3.2 3B, Llama‑3.2 1B. Blue traces denote training loss, red traces denote validation loss; the vertical dashed line indicates the checkpoint selected for evaluation.
IV Performance and Evaluation
IV-A Evaluation metrics and protocol conformance
[…]
Semantic similarity (SBERT cosine)
For each test sample $i$, let $Y_{i}$ be the ground-truth DL message and $\hat{Y}_{i}$ the model output given the UL context $X_{i}$. We encode both with a frozen SBERT encoder $f_{\text{sbert}}(\cdot)$ and mean-pool to sentence embeddings $u_{i}$ and $v_{i}$:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u_{i}=\frac{1}{|Y_{i}|}\sum_{t}f_{\text{sbert}}(Y_{i})_{t},\qquad v_{i}=\frac{1}{|\hat{Y}_{i}|}\sum_{t}f_{\text{sbert}}(\hat{Y}_{i})_{t}.$} (5)
The cosine similarity is $s_{i}=\frac{u_{i}^{\top}v_{i}}{\|u_{i}\|\,\|v_{i}\|}\in[-1,1]$. We report $s_{i}$ directly. When combining semantic and structural metrics, a $[0,1]$ range is preferred, so we use $\tilde{s}_{i}=\frac{s_{i}+1}{2}$. Fig. 7 illustrates the distribution of SBERT-based similarity scores for the 5G NR field corpus under the LLaMA-3 8B backbone after fine-tuning with LoRA.
ASN.1 round-trip validity and schema recall
For each session $i$, we construct a trimmed ASN.1 micro-schema $\mathcal{S}^{(i)}$ (messages and IEs relevant to that session) and perform a strict encode–decode round-trip on $\hat{Y}_{i}$. The binary pass indicator is
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{asn}_{i}\;=\;\mathbb{I}\Big[\mathrm{Dec}_{\mathcal{S}^{(i)}}\big(\mathrm{Enc}_{\mathcal{S}^{(i)}}(\hat{Y}_{i})\big)\ \text{succeeds $\land$ plausible}\Big].$} (6)
To measure field-level faithfulness beyond mere text similarity, we compare the set of normalized ASN.1 Information Element (IE) names present in the reference and the hypothesis. Let $W(\cdot)$ extract this set; for non-empty $W(Y_{i})$, the per-sample schema recall is
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r_{i}\;=\;\frac{\big|\,W(Y_{i})\cap W(\hat{Y}_{i})\,\big|}{\big|\,W(Y_{i})\,\big|}\in[0,1].$} (7)
Figure 7: Distribution of cosine similarity scores (simple pooling-based metric, not SBERT) on the NR dataset (proprietary corpus) using the LLaMA-3 8B backbone. The histogram contrasts the fine-tuned model against zero-shot baselines.
[…]
From the UL input we derive the set $U_{i}$ of UL message types; from $\hat{Y}_{i}$ we extract the DL type $d_{i}$ and, when present, the RRC transaction identifier $\mathrm{TxId}(\hat{Y}_{i})$. Let $\mathcal{A}(u)$ denote the set of allowed DL types given a UL type $u$ (strong edges); for a set $U$, write $\mathcal{A}(U)=\bigcup_{u\in U}\mathcal{A}(u)$. The strict binary SMC pass is then
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{smc}_{i}=\mathbb{I}\!\Big[d_{i}\in\mathcal{A}(U_{i})\ \wedge\ \mathrm{TxId}(\hat{Y}_{i})=\mathrm{TxId}(Y_{i})\Big].$} (8)
Here, the transaction-ID equality is enforced only when both sides carry an ID. We additionally report a weak profile in which broadcast/weak edges are admitted (not shown for brevity when identical).
Reporting
In Tables IVV, we summarize results by the distribution (median and dispersion) of $s_{i}$ (or $\tilde{s}_{i}$), the dataset-level ASN.1 pass rate $\frac{1}{N}\sum_{i}\mathrm{asn}_{i}$, the schema-recall distribution of $r_{i}$, and the SMC pass rate $\frac{1}{N}\sum_{i}\mathrm{smc}_{i}$. These metrics are presented side-by-side with latency to expose fidelity–validity–latency trade-offs without conflating them into a single score.
TABLE V: Global medians by decoding regime and precision across all 120 configurations on the LTE dataset
Regime Precision Latency (ms) Similarity (med.) Schema (med.) ASN (med.) SMC (med.)
NoSys FP16 1982 0.506 0.208 0.008 0.004
NoSys Q4_K_M 1480 0.502 0.209 0.013 0.004
RRC FP16 1697 0.606 0.267 0.003 0.001
RRC Q4_K_M 1277 0.579 0.238 0.004 0.001
RRC_constrain FP16 2338 1.000 1.000 0.994 0.991
RRC_constrain Q4_K_M 1889 1.000 0.963 0.835 0.831
IV-B Inference results: fidelity, validity, and latency
This subsection reports empirical outcomes for semantic fidelity (SBERT cosine), ASN.1 validity, SMC pass rate, and latency across the evaluated configurations. We focus on the curated highlights (Table IV) and global medians over all 120 settings (Table V).
TABLE VI: Per‑procedure metrics on the LTE dataset
Procedure SMC pass ASN1 true Sim. Schema
Information Transfer 1.000 1.000 0.969 0.895
Initial Security Activation 1.000 1.000 0.988 0.989
Intra-EUTRA Handover 0.875 0.875 0.700 0.457
RRC Connection Establishment 0.995 0.999 0.994 0.980
RRC Connection Re-establishment 1.000 1.000 0.987 0.987
RRC Connection Reconfiguration 0.988 0.992 0.987 0.951
RRC Connection Release 0.990 1.000 0.974 0.943
UE Capability Transfer 0.997 0.997 0.996 0.995
Constrained decoding is decisive
Under RRC_constrain, fine-tuned backbones (8B/3B/1B) achieve near-ceiling ASN.1/SMC pass (typically 0.971.000.97\!-\!1.00) with saturated semantic similarity, whereas the same models with a generic prompt (RRC) exhibit high textual similarity yet near-zero ASN.1/SMC in many cases (Table IV). This confirms that schema-bounded prompting is necessary to convert semantic proximity into standards-compliant messages.
TABLE VII: Cross‑operator metrics on the LTE dataset
Operator SMC pass ASN1 true Sim. Schema
Operator A 0.549 0.997 0.975 0.931
Operator B 0.536 0.999 0.969 0.913
Parameter-efficient tuning vs. full fine-tuning
With RRC_constrain, LoRA at rank r=16r{=}16 matches or nearly matches full fine-tuning on 8B/3B models in both similarity and validator success. Smaller backbones (1B) generally favor full updates for peak validity (Table IV). Thus, parameter-efficient adaptation suffices at larger scales, while compact models benefit from full FT.
Quantization–latency trade-off (weight-only INT4)
INT4 (Q4_K_M) consistently reduces median latency by $\approx\!20\mbox{--}30\%$ relative to FP16 at comparable decoding settings, with small average reductions in ASN.1/SMC pass (typically a few percentage points and backbone-dependent). The largest drops appear on the smallest backbone (1B), consistent with reduced numerical headroom (Table IV, Table V).
Latency snapshot
Under RRC_constrain + Q4_K_M, origin 3B reaches median latencies in the few-hundred-millisecond range ($\sim 280\mbox{--}370$ ms) at reduced validator pass. In contrast, fine-tuned models retain near-ceiling ASN.1/SMC at medians $\sim 1.7\mbox{--}3.6$ s depending on backbone and tuning (Table IV, Table V).
Per-procedure and cross-operator behavior
The Long Term Evolution (LTE) benchmark shows near-ceiling pass rates for most procedures, with Intra-Evolved UMTS Terrestrial Radio Access (E-UTRA) handover as the primary bottleneck (Table VI). Anonymized cross-operator results indicate consistent validity across deployments (Table VII).
Zero-shot public LAM baselines (context)
On the field-collected subset, the specialized RRC-LLM attains a much higher median semantic similarity than zero-shot public models (Table VIII). This contextualizes the gain from domain adaptation.
TABLE VIII: Median cosine similarity of the fully fine-tuned LLaMA 8B backbone
Model NR dataset
Gemini 1.5‑Flash 0.585
GPT‑4o 0.496
Claude 3.5 Sonnet v2 0.768
GPT-o3‑mini 0.705
Original Llama 3‑8B 0.600
RRC‑LLM (Ours) 0.970
TABLE IX: Examples of RRC-LLM Responses Across Similarity Levels
High-Similarity (Similarity = 0.9991)
True Label: RRC Message RRC-LLM (Similarity = 0.9991)
Time: 2024-04-15 11:34:38.469892 Message name: rrcSetup Interface: F1_E1 Protocol: nrRrc Time diff: 0000:00:00.001898 Message body: DL-CCCH-Message: { message c1 : rrcSetup : { rrc-TransactionIdentifier 0, criticalExtensions rrcSetup: { radioBearerConfig { srb-ToAddModList { { srb-Identity 1 } } ... Time: 2024-04-15 11:34:38.469982 Message name: rrcSetup Interface: F1_E1 Protocol: nrRrc Time diff: 0000:00:00.001988 Message body: DL-CCCH-Message: { message c1 : rrcSetup : { rrc-TransactionIdentifier 0, criticalExtensions rrcSetup: { radioBearerConfig { srb-ToAddModList { { srb-Identity 1 } } ...
Medium-Similarity (Similarity = 0.92)
True Label: RRC Message RRC-LLM (Similarity = 0.92)
... headerCompression notUsed : NULL, statusReportRequired true }, t-Reordering ms100 } } } }, measConfig { measObjectToAddModList { { measObjectId 1, measObject measObjectNR : { ssbFrequency 634080, ssbSubcarrierSpacing kHz30, smtc1 { periodicityAndOffset sf20 : 0, duration sf5 }, ... ... headerCompression notUsed : NULL }, t-Reordering ms100 } } }, securityConfig { securityAlgorithmConfig { cipheringAlgorithm nea2, integrityProtAlgorithm nia2 } } }, measConfig { measObjectToAddModList { { measObjectId 1, measObject measObjectNR : { ssbFrequency 634080, ssbSubcarrierSpacing kHz30, smtc1 { periodicityAndOffset sf20 : 0, duration sf5 }, ...
Low-Similarity (Similarity = 0.68)
True Label: RRC Message RRC-LLM (Similarity = 0.68)
Time: 2024-04-15 11:33:10.690388 Message name: rrcRelease Interface: F1_E1 Protocol: nrRrc Time diff: 0000:00:10.633832 Message body: DL-DCCH-Message: { message c1 : rrcRelease : { rrc-TransactionIdentifier 0, criticalExtensions rrcRelease: { } } } ... It looks like theres another RRC reconfiguration Complete message being sent! Message name: rrcReconfigurationComplete Time: 2024-04-15 11:33:00.243997 Interface: F1_E1 Protocol: nrRrc Time diff: 0000:00:00.187441 Message body: UL-DCCH-Message: { message c1 : rrcReconfigurationComplete : { rrc-TransactionIdentifier 3, criticalExtensions rrcReconfigurationComplete: { } } } ...
Takeaway
Across 120 configurations, the main conclusions are: (i) RRC_constrain is required for standards conformance; (ii) large backbones with LoRA r=16r{=}16 match full FT under constraints, while 1B favors full FT; (iii) INT4 delivers substantial latency gains at modest validity cost; and (iv) the tuned, standards-compliant configurations remain above sub-100 ms median latency, leaving further acceleration to future engineering (Sec. V).
TABLE X: Llama-3.2-3B at Q4_K_M under RRC_constrain: latency vs. fidelity vs. validity (medians) on the LTE dataset
Tuning Latency (ms) Similarity Schema ASN SMC
origin 280 0.601 0.286 0.390 0.046
LoRA r=4 1723 0.999 0.941 0.748 0.743
LoRA r=8 1681 0.999 0.948 0.671 0.666
LoRA r=16 1730 1.000 1.000 0.985 0.984
full FT 1733 1.000 1.000 0.973 0.968
IV-C Latency and inference throughput
Latency is reported as the median time to generate a single DL RRC message under the evaluated decoding regimes and precision settings. All measurements were taken on high-performance accelerators described in Table XI, using identical test splits as in Tables IVV.
TABLE XI: Hardware platforms used for training and inference
Component Specification
Used on NR dataset
CPU 52-core hyper-threaded
RAM 512 GB
GPU 2 $\times$ NVIDIA A100 80 GB (PCIe)
Interconnect PCIe peer-to-peer (GPU–GPU), PCIe switch (GPU–host)
GPU VRAM (Inference) $\sim$29 GB per thread (GGUF format)
GPU VRAM (Training) $\sim$110 GB across two cards
CUDA Version v12
Used on LTE dataset
Cluster Size 1,320 nodes
Node Composition $4\times$ NVIDIA GH200 (Grace 72-core CPU + H100 96 GB HBM3)
Per-Node Memory Grace: $4\times 120$ GB LPDDR5X; H100: $4\times 96$ GB HBM3; $\approx$844 GB usable
[…]
CPU–GPU Link NVLink-C2C 900 GB/s (coherent)
GPU–GPU (Intra-node) NVLink/NVSwitch
Inter-node Network HPE Cray Slingshot 11 ($4\times 200$ Gb/s per node)
Power Caps $\sim$660 W per GH200 (CPU+GPU); $\sim$2.64 kW per node
CUDA Version v12
Edge Device
Platform Apple M2 Max (12-core CPU + 30-core GPU)
Memory 32 GB unified memory
Unified Memory Bandwidth $\sim$400 GB/s
Estimated Peak Compute $\sim$10.65 TFLOPS FP32 (third-party)
Observed ranges
Under the generic prompt (RRC), fine-tuned 8B backbones exhibit medians exceeding 3 s, with tails beyond 5 s. Introducing schema-bounded decoding (RRC_constrain) reduces variance and stabilizes output length, but tuned configurations remain in the 1.7–3.6 s range depending on backbone and tuning method (Table IV).
Figure 8: Latency–fidelity trade-off across configurations on the LTE dataset. Each point represents a fine-tuning/quantization/backbone setting. Color encodes decoding regime, marker shape encodes quantization, size encodes backbone scale, and fill intensity encodes fine-tuning depth. Vertical dashed lines mark RRC timers T304 ($512$ ms) and T300 ($1000$ ms).
Effect of quantization
Weight-only INT4 (Q4_K_M) lowers medians by approximately 20–30% relative to FP16 without altering the decoding regime. For example, an 8B model under RRC_constrain drops from $\sim$3.6 s to $\sim$2.6 s, while a 3B backbone falls from $\sim$2.3 s to $\sim$1.7 s (Table V). Validator success remains high, with only minor reductions in ASN.1/SMC pass rates. Figure 8 visualizes the latency–fidelity frontier for all LTE configurations. The fastest points correspond to origin models, which achieve low latency by producing short, often invalid RRC messages (reflected in their poor semantic similarity). Conversely, tuned configurations deliver near-ceiling protocol validity (Tables VI–VII) but remain well above RRC timer budgets, with the best case at $\sim$1.6 s. Notably, smaller backbones (1B/3B) reduce latency substantially without compromising semantic fidelity, suggesting that backbone scaling combined with quantization and constrained decoding is a promising direction for future sub-100 ms targets.
Lower-bound configurations
Configurations labeled as origin refer to base models without any fine-tuning. When combined with RRC_constrain and INT4, these achieve the shortest medians observed ($\sim$280–370 ms for 3B), but do not meet the fidelity and validity thresholds required for deployment. This illustrates the latency–conformance trade-off inherent in current designs.
On the sub-100 ms target
No tuned configuration achieves sub-100 ms median latency. Meeting this requirement will demand additional engineering measures such as speculative multi-token decoding, hierarchical agents, and lightweight backbones.
Edge–vs.–datacenter under 1B/INT4 (M2 vs. GH200)
[…]
TABLE XII: Edge portability with Llama-3.2-1B / INT4 (Q4_K_M) under RRC_constrain. Medians read from the M2 vs. GH200 distributions; energy is a conservative upper bound using platform power caps.
Platform Median Lat. (ms) Throughput (req/s) Schema chk (med) Sim. (med) SMC conf. Energy/ msg
GH200 ($\timesH1$$0$) $\approx$1092 $\approx$0.92 $\approx$0.96 1.00 0.53 $\approx$0.200 Wh
Apple M2 Max $\approx$1841 $\approx$0.54 $\approx$0.97 1.00 0.54 $\approx$7.7 mWh
V Conclusion and Future Work
This paper demonstrated the feasibility of embedding a decoder-only LAM within the RRC layer of a disaggregated gNB, achieving standards-compliant message generation through supervised fine-tuning and schema-bounded prompting. Across 120 configurations, the proposed approach attains near-ceiling ASN.1 validity and SMC pass rates under RRC_constrain, while parameter-efficient tuning (LoRA) approaches full fine-tuning performance on large backbones. Quantization (Q4_K_M) delivers latency reductions of 20–30% without compromising structural fidelity.
Despite these gains, two limitations remain: (i) tuned configurations do not meet sub-100 ms median latency targets, and (ii) current evaluations rely on high-performance accelerators (Table XI). Telecom-grade hardware validation is deferred.
Future work will pursue four directions:
  1. 1.
    Inference acceleration: speculative multi-token decoding [7], hierarchical agent architectures, and lightweight backbones to reconcile fidelity with real-time constraints.
  2. 2.
    Extended-context modeling: the context window length of current LLM s is critical for cross-layer reasoning and complex RRC procedures. Longer contexts or memory-augmented transformers (e.g., RoPE-extended variants, ALiBi [11]) will be explored to eliminate segmentation and preserve procedural continuity.
  3. 3.
    Dynamic schema adaptation: retrieval-augmented generation for evolving 3GPP releases and vendor-specific variants.
  4. 4.
    Protocol-aware evaluation metrics: current metrics (Exact Match, SBERT similarity) are insufficient. Exact Match penalizes benign variations (e.g., transaction identifiers, minor threshold offsets), while SBERT ignores protocol semantics and network-level KPIs (e.g., RRC Setup Rejects, failed handovers). Future work will define an RRC-specific protocol score that integrates structural validity, semantic intent, and operational KPIs for a more meaningful assessment of control-plane performance.
We believe the research trajectory outlined here closes the loop between the conceptual promise of an AI-AI and its concrete instantiation in next-generation RAN deployments.
Appendix A Dataset Construction and Prompt Templates
TABLE A.1: Symbols used in LTE sessionization and prompt construction.
Symbol Meaning
$K$ Q/A turns actually produced for a session.
$K_{\max}$ Max. Q/A turns per LTE session (turn cap).
$\kappa_{\mathrm{SIB}},\ \kappa_{\mathrm{PG}}$ Max retained SIB1/Paging per session (merged into DL, never standalone).
$\sigma$ Intra-block join delimiter for adjacent concatenation inside $Q_{k}$$A_{k}$.
$L_{\max}$ Context/window budget for a serialized session (tokens).
$B_{\mathrm{ASN1}}$ Token budget of the session-specific LTE ASN.1 micro-schema injected in system.
$W_{\mathrm{start}}$ Window anchored at the session start for broadcast retention.
$W_{\mathrm{pre}}$ Prefix window preceding the target DL block when merging broadcasts.
$g_{\mathrm{thin}}(\cdot;\Theta_{\mathrm{thin}})$ Optional thinning of repeated MeasurementReport lines within the same $Q_{k}$.
$C^{(\mathrm{blk})}_{\max}$ Per-block serialized size bound for any single $Q_{k}$$A_{k}$ (tokens).
$\mathrm{PCI}_{\mathrm{allowed}}$ Allowed physical-cell IDs (serving, targets in mobilityControlInfo, frequent neighbors).
$\mathrm{EARFCN}_{\mathrm{allowed}}$ Allowed EARFCN set (serving carrier and any targets referenced in session).
[…]
(a) Canonical message labels
Message(s) Role / channel (illustrative)
RRCConnectionRequest, RRCConnectionSetup RRC connection establishment anchors (UL/DL-CCCH)
RRCConnectionSetupComplete, SecurityModeCommand/SecurityModeComplete Security start (DL/UL-DCCH)
RRCConnectionReconfiguration/RRCConnectionReconfigurationComplete (Re)configuration (DL/UL-DCCH)
MeasurementReport Mobility trigger (UL-DCCH)
ULInformationTransfer/DLInformationTransfer NAS transport (DCCH)
RRCConnectionReestablishmentRequest/RRCConnectionReestablishment/RRCConnectionReestablishmentComplete Reestablishment (UL-CCCH / DL-CCCH / UL-DCCH)
RRCConnectionResumeRequest/RRCConnectionResume/RRCConnectionResumeComplete Resume (UL-CCCH / DL-CCCH / UL-DCCH)
RRCConnectionRelease Session end (DL-DCCH)
SystemInformationBlockType1(SIB1) BCCH-DL-SCH (broadcast; retained $\leq 1$; no PCI in SIB1)
Paging PCCH (broadcast; retained $\leq 1$)
[…]
UL message(s) Typical DL reply (notes)
RRCConnectionRequest (UL-CCCH) RRCConnectionSetup (DL-CCCH)
RRCConnectionSetupComplete (UL-DCCH) SecurityModeCommand (DL-DCCH), then SecurityModeComplete (UL-DCCH), then RRCConnectionReconfiguration (DL-DCCH)
MeasurementReport (UL-DCCH) RRCConnectionReconfiguration (often with mobilityControlInfo)
ULInformationTransfer (UL-DCCH, NAS) DLInformationTransfer (DL-DCCH, NAS) or start of security/configuration
UECapabilityInformation (UL-DCCH) RRCConnectionReconfiguration (DL-DCCH)
RRCConnectionReestablishmentRequest (UL-CCCH) RRCConnectionReestablishment (DL-CCCH) $\rightarrow$ RRCConnectionReestablishmentComplete (UL-DCCH)
RRCConnectionResumeRequest (UL-CCCH) RRCConnectionResume (DL-CCCH) $\rightarrow$ RRCConnectionResumeComplete (UL-DCCH)
RRCConnectionReconfigurationComplete (UL-DCCH) Often none; or policy-dependent RRCConnectionRelease (DL-DCCH)
Figure A.1: Session-specific system prompt used to seed the conversation.
Appendix B Full Benchmark Results
TABLE B.1: Full results on the LTE dataset (complete results across backbones, quantizations, strategies, and tuning).
Backbone Quantization Decoding regime Fine Tuning Latency (ms) Schema Similarity Pass Rate
Med. Avg. Avg Med Avg Med ASN SMC
L-3 8B FP16 RRC_constrain full 3597 5330.5 0.967 1.000 0.988 1.000 0.997 0.994
L-3 8B FP16 RRC_constrain lora-r16 3544 5258.0 0.976 1.000 0.992 1.000 0.995 0.993
L-3 8B FP16 RRC_constrain lora-r8 3571 5297.5 0.971 1.000 0.990 1.000 0.995 0.992
L-3 8B FP16 RRC_constrain lora-r4 3608 5353.7 0.972 1.000 0.992 1.000 0.997 0.996
L-3 8B FP16 RRC_constrain origin 394 549.4 0.715 0.833 0.732 0.738 0.091 0.088
L-3 8B FP16 RRC full 3522 4221.4 0.464 0.444 0.694 0.695 0.000 0.000
L-3 8B FP16 RRC lora-r16 3418 4280.4 0.467 0.333 0.723 0.669 0.023 0.000
L-3 8B FP16 RRC lora-r8 3562 5403.3 0.657 0.881 0.786 0.760 0.003 0.001
L-3 8B FP16 RRC lora-r4 738 3916.2 0.296 0.265 0.610 0.633 0.051 0.047
L-3 8B FP16 RRC origin 1095 1131.8 0.237 0.242 0.498 0.520 0.000 0.000
L-3 8B FP16 NoSys full 3892 4694.6 0.419 0.365 0.657 0.652 0.001 0.000
L-3 8B FP16 NoSys lora-r16 1987 2011.7 0.231 0.241 0.480 0.478 0.003 0.002
L-3 8B FP16 NoSys lora-r8 2177 2198.6 0.226 0.243 0.504 0.506 0.000 0.000
L-3 8B FP16 NoSys lora-r4 2082 2200.1 0.222 0.226 0.509 0.509 0.001 0.000
L-3 8B FP16 NoSys origin 2018 2098.2 0.236 0.228 0.497 0.490 0.003 0.002
L-3 8B Q4_K_M RRC_constrain full 2584 3903.8 0.966 1.000 0.989 1.000 0.998 0.993
L-3 8B Q4_K_M RRC_constrain lora-r16 2556 3854.1 0.969 1.000 0.990 1.000 0.989 0.988
L-3 8B Q4_K_M RRC_constrain lora-r8 2562 3903.4 0.961 1.000 0.989 1.000 0.942 0.940
L-3 8B Q4_K_M RRC_constrain lora-r4 2571 3855.1 0.941 0.978 0.986 1.000 0.899 0.897
L-3 8B Q4_K_M RRC_constrain origin 368 412.8 0.597 0.625 0.709 0.728 0.152 0.147
L-3 8B Q4_K_M RRC full 2407 3223.6 0.392 0.357 0.687 0.691 0.004 0.001
L-3 8B Q4_K_M RRC lora-r16 2519 2529.7 0.439 0.333 0.726 0.666 0.003 0.002
L-3 8B Q4_K_M RRC lora-r8 2486 3820.7 0.560 0.658 0.755 0.724 0.007 0.001
L-3 8B Q4_K_M RRC lora-r4 1331 1867.2 0.405 0.333 0.625 0.628 0.002 0.001
L-3 8B Q4_K_M RRC origin 900 958.2 0.220 0.212 0.546 0.553 0.001 0.001
L-3 8B Q4_K_M NoSys full 2154 2537.2 0.443 0.444 0.676 0.680 0.003 0.000
L-3 8B Q4_K_M NoSys lora-r16 1431 1442.1 0.235 0.239 0.491 0.490 0.012 0.004
L-3 8B Q4_K_M NoSys lora-r8 1447 1529.0 0.229 0.240 0.503 0.500 0.000 0.000
L-3 8B Q4_K_M NoSys lora-r4 1570 1627.8 0.228 0.236 0.510 0.505 0.001 0.001
L-3 8B Q4_K_M NoSys origin 1435 1531.4 0.233 0.239 0.514 0.508 0.002 0.001
L-3.1 8B FP16 RRC_constrain full 3631 5262.4 0.973 1.000 0.990 1.000 0.997 0.994
L-3.1 8B FP16 RRC_constrain lora-r16 3609 5367.3 0.969 1.000 0.991 1.000 0.995 0.994
L-3.1 8B FP16 RRC_constrain lora-r8 3587 5309.9 0.970 1.000 0.990 1.000 0.996 0.994
L-3.1 8B FP16 RRC_constrain lora-r4 3591 5308.4 0.967 1.000 0.989 1.000 0.998 0.997
L-3.1 8B FP16 RRC_constrain origin 840 1075.5 0.393 0.368 0.678 0.711 0.131 0.118
L-3.1 8B FP16 RRC full 3324 2625.7 0.372 0.417 0.712 0.712 0.005 0.001
L-3.1 8B FP16 RRC lora-r16 3364 2716.1 0.523 0.675 0.643 0.650 0.001 0.000
L-3.1 8B FP16 RRC lora-r8 1743 2195.3 0.382 0.275 0.575 0.520 0.002 0.001
L-3.1 8B FP16 RRC lora-r4 1997 2461.2 0.249 0.218 0.545 0.557 0.002 0.000
L-3.1 8B FP16 RRC origin 2153 2261.3 0.208 0.220 0.520 0.524 0.001 0.001
L-3.1 8B FP16 NoSys full 1981 2112.1 0.234 0.216 0.518 0.513 0.025 0.010
L-3.1 8B FP16 NoSys lora-r16 2193 2243.0 0.211 0.224 0.477 0.478 0.003 0.002
L-3.1 8B FP16 NoSys lora-r8 2119 2288.3 0.225 0.215 0.511 0.504 0.002 0.001
L-3.1 8B FP16 NoSys lora-r4 2686 2766.8 0.202 0.194 0.501 0.504 0.033 0.012
L-3.1 8B FP16 NoSys origin 2791 2837.0 0.183 0.182 0.478 0.478 0.006 0.003
L-3.1 8B Q4_K_M RRC_constrain full 2571 3798.8 0.973 1.000 0.990 1.000 0.994 0.993
L-3.1 8B Q4_K_M RRC_constrain lora-r16 2581 3939.0 0.964 0.987 0.989 1.000 0.904 0.902
L-3.1 8B Q4_K_M RRC_constrain lora-r8 2569 3858.3 0.933 0.987 0.983 1.000 0.965 0.962
L-3.1 8B Q4_K_M RRC_constrain lora-r4 2547 4079.2 0.905 0.923 0.987 0.998 0.653 0.651
L-3.1 8B Q4_K_M RRC_constrain origin 481 674.0 0.394 0.385 0.662 0.674 0.138 0.099
L-3.1 8B Q4_K_M RRC full 2420 2349.7 0.470 0.417 0.710 0.679 0.010 0.001
L-3.1 8B Q4_K_M RRC lora-r16 1379 1653.6 0.371 0.262 0.579 0.508 0.001 0.000
L-3.1 8B Q4_K_M RRC lora-r8 1064 1218.5 0.260 0.241 0.509 0.491 0.002 0.002
L-3.1 8B Q4_K_M RRC lora-r4 1360 1559.6 0.286 0.255 0.570 0.554 0.007 0.003
L-3.1 8B Q4_K_M RRC origin 1576 1617.9 0.212 0.219 0.517 0.506 0.001 0.000
L-3.1 8B Q4_K_M NoSys full 1469 1470.3 0.230 0.225 0.566 0.568 0.418 0.138
L-3.1 8B Q4_K_M NoSys lora-r16 1501 1537.8 0.213 0.220 0.485 0.482 0.001 0.000
L-3.1 8B Q4_K_M NoSys lora-r8 1674 1753.3 0.215 0.212 0.484 0.480 0.001 0.000
L-3.1 8B Q4_K_M NoSys lora-r4 1872 1878.7 0.208 0.202 0.499 0.501 0.014 0.003
L-3.1 8B Q4_K_M NoSys origin 1886 1931.0 0.197 0.196 0.491 0.484 0.001 0.001
L-3.2 1B FP16 RRC_constrain full 2093 3867.2 0.964 1.000 0.988 1.000 0.993 0.989
L-3.2 1B FP16 RRC_constrain lora-r16 2048 3775.5 0.964 1.000 0.988 1.000 0.993 0.989
L-3.2 1B FP16 RRC_constrain lora-r8 2146 3928.3 0.955 0.993 0.986 1.000 0.988 0.985
L-3.2 1B FP16 RRC_constrain lora-r4 2162 3963.2 0.961 1.000 0.989 1.000 0.992 0.989
L-3.2 1B FP16 RRC_constrain origin 338 472.1 0.274 0.250 0.632 0.679 0.215 0.112
L-3.2 1B FP16 RRC full 589 1251.0 0.336 0.333 0.652 0.649 0.000 0.000
L-3.2 1B FP16 RRC lora-r16 959 1421.7 0.377 0.312 0.650 0.660 0.002 0.002
L-3.2 1B FP16 RRC lora-r8 1165 2331.0 0.258 0.200 0.596 0.583 0.004 0.002
L-3.2 1B FP16 RRC lora-r4 1960 2922.9 0.343 0.270 0.643 0.630 0.004 0.001
L-3.2 1B FP16 RRC origin 1075 1123.4 0.234 0.243 0.551 0.551 0.013 0.004
L-3.2 1B FP16 NoSys full 1142 1390.6 0.202 0.200 0.521 0.514 0.024 0.009
L-3.2 1B FP16 NoSys lora-r16 1704 2143.1 0.198 0.187 0.504 0.505 0.008 0.004
L-3.2 1B FP16 NoSys lora-r8 1562 2230.3 0.199 0.190 0.503 0.504 0.016 0.007
L-3.2 1B FP16 NoSys lora-r4 1462 1786.2 0.192 0.193 0.515 0.512 0.008 0.005
L-3.2 1B FP16 NoSys origin 1281 1360.5 0.203 0.210 0.527 0.520 0.007 0.002
L-3.2 1B Q4_K_M RRC_constrain full 2025 3585.8 0.934 0.987 0.981 1.000 0.888 0.883
L-3.2 1B Q4_K_M RRC_constrain lora-r16 1878 3581.9 0.829 0.923 0.969 0.998 0.548 0.542
L-3.2 1B Q4_K_M RRC_constrain lora-r8 1880 3455.9 0.690 0.736 0.960 0.995 0.516 0.508
L-3.2 1B Q4_K_M RRC_constrain lora-r4 1897 3512.5 0.854 0.947 0.973 0.998 0.783 0.779
L-3.2 1B Q4_K_M RRC_constrain origin 339 769.1 0.225 0.217 0.642 0.656 0.215 0.095
L-3.2 1B Q4_K_M RRC full 428 1368.8 0.293 0.286 0.662 0.665 0.000 0.000
L-3.2 1B Q4_K_M RRC lora-r16 513 1566.0 0.244 0.221 0.600 0.600 0.003 0.000
L-3.2 1B Q4_K_M RRC lora-r8 1202 2794.3 0.208 0.188 0.626 0.615 0.008 0.001
L-3.2 1B Q4_K_M RRC lora-r4 1874 2568.6 0.279 0.235 0.630 0.627 0.005 0.002
L-3.2 1B Q4_K_M RRC origin 1118 1208.6 0.213 0.215 0.526 0.525 0.006 0.003
L-3.2 1B Q4_K_M NoSys full 1209 1335.5 0.189 0.182 0.501 0.502 0.024 0.016
L-3.2 1B Q4_K_M NoSys lora-r16 1602 1982.3 0.183 0.181 0.494 0.494 0.004 0.002
L-3.2 1B Q4_K_M NoSys lora-r8 1825 2689.0 0.187 0.183 0.499 0.501 0.013 0.008
L-3.2 1B Q4_K_M NoSys lora-r4 1305 1567.4 0.176 0.172 0.496 0.502 0.008 0.004
L-3.2 1B Q4_K_M NoSys origin 1417 1514.2 0.203 0.202 0.516 0.515 0.050 0.016
L-3.2 3B FP16 RRC_constrain full 2341 3501.9 0.969 1.000 0.991 1.000 0.993 0.991
L-3.2 3B FP16 RRC_constrain lora-r16 2341 3502.4 0.972 1.000 0.991 1.000 0.998 0.995
L-3.2 3B FP16 RRC_constrain lora-r8 2335 3533.8 0.966 1.000 0.989 1.000 0.994 0.991
L-3.2 3B FP16 RRC_constrain lora-r4 2331 3492.7 0.963 1.000 0.989 1.000 0.994 0.992
L-3.2 3B FP16 RRC_constrain origin 369 360.5 0.369 0.400 0.549 0.548 0.328 0.035
L-3.2 3B FP16 RRC full 2157 2939.1 0.408 0.353 0.688 0.700 0.000 0.000
L-3.2 3B FP16 RRC lora-r16 1467 1930.6 0.233 0.229 0.555 0.549 0.003 0.001
L-3.2 3B FP16 RRC lora-r8 1566 1645.1 0.206 0.209 0.556 0.562 0.031 0.010
L-3.2 3B FP16 RRC lora-r4 1651 1647.4 0.218 0.221 0.527 0.524 0.035 0.013
L-3.2 3B FP16 RRC origin 1645 1676.8 0.221 0.220 0.550 0.548 0.023 0.009
L-3.2 3B FP16 NoSys full 1755 1848.0 0.214 0.223 0.466 0.467 0.011 0.004
L-3.2 3B FP16 NoSys lora-r16 1894 1906.7 0.196 0.200 0.524 0.537 0.034 0.012
L-3.2 3B FP16 NoSys lora-r8 1962 2021.1 0.203 0.207 0.524 0.520 0.116 0.034
L-3.2 3B FP16 NoSys lora-r4 1983 2084.9 0.206 0.204 0.518 0.518 0.154 0.091
L-3.2 3B FP16 NoSys origin 1946 2005.4 0.202 0.205 0.499 0.499 0.015 0.008
L-3.2 3B Q4_K_M RRC_constrain full 1733 2658.1 0.954 1.000 0.987 1.000 0.973 0.968
L-3.2 3B Q4_K_M RRC_constrain lora-r16 1730 2629.4 0.969 1.000 0.989 1.000 0.985 0.984
L-3.2 3B Q4_K_M RRC_constrain lora-r8 1681 2628.2 0.914 0.948 0.982 0.999 0.671 0.666
L-3.2 3B Q4_K_M RRC_constrain lora-r4 1723 2562.1 0.867 0.941 0.976 0.999 0.748 0.743
L-3.2 3B Q4_K_M RRC_constrain origin 280 302.4 0.306 0.286 0.590 0.601 0.390 0.046
L-3.2 3B Q4_K_M RRC full 2170 2920.6 0.440 0.400 0.689 0.698 0.001 0.000
L-3.2 3B Q4_K_M RRC lora-r16 1105 1117.6 0.212 0.222 0.524 0.528 0.005 0.002
L-3.2 3B Q4_K_M RRC lora-r8 1224 1264.1 0.211 0.210 0.571 0.558 0.052 0.022
L-3.2 3B Q4_K_M RRC lora-r4 1192 1223.1 0.216 0.215 0.536 0.533 0.004 0.001
L-3.2 3B Q4_K_M RRC origin 1105 1156.1 0.215 0.212 0.562 0.557 0.061 0.026
L-3.2 3B Q4_K_M NoSys full 1367 1452.8 0.201 0.207 0.497 0.493 0.040 0.015
L-3.2 3B Q4_K_M NoSys lora-r16 1328 1387.4 0.202 0.216 0.548 0.543 0.033 0.020
L-3.2 3B Q4_K_M NoSys lora-r8 1518 1556.5 0.189 0.197 0.514 0.514 0.058 0.022
L-3.2 3B Q4_K_M NoSys lora-r4 1491 1517.3 0.206 0.207 0.513 0.512 0.164 0.098
L-3.2 3B Q4_K_M NoSys origin 1371 1402.1 0.209 0.211 0.520 0.515 0.026 0.011